# The Structure of the Majorana Clifford Group

### v2

@@ A.2 Theorem 4
where the second line is a direct consequence of
$\sum_{v\in M}(-1)^{a^{T}v}=|M|\,\delta_{a\in M^{\perp}}$ (102)
for any subspace $M\subseteq\mathbb{F}_{2}^{2n}$ and $M^{\perp}=\{v\in\mathbb{F}_{2}^{2n}\,|\,v^{T}v^{\prime}=0\quad\forall v^{\prime}\in M\}$. This also implies that
$\sum_{a\in\mathbb{F}_{2}^{2n,+}}\lvert\operatorname{tr}[\Pi_{+}\operatorname{\gamma}(S)\operatorname{\mu}(a)]\rvert^{2}=2^{2n-1},$ (103)
since $(\mathbb{F}_{2}^{2n,+})^{\perp}\cap\mathbb{F}_{2}^{2n,+}=\{0,j\}$, and both elements are fixed points. Hence in this sum $2^{2n}/(f_{+}(S)+c_{+}(S))$ terms are non-zero and therefore equal to (101). It follows that
$\begin{split}&\sum_{a\in\mathbb{F}_{2}^{2n,+}}\lvert\operatorname{tr}[\Pi_{+}\operatorname{\gamma}(S)\operatorname{\mu}(a)]\rvert^{2t}\\ ={}&\frac{2^{2n}}{(f_{+}(S)+c_{+}(S))}\cdot\left[\frac{f_{+}(S)+c_{+}(S)}{2}\right]^{t}\\ ={}&\lvert\mathbb{F}_{2}^{2n,+}\rvert\cdot\left[\frac{f_{+}(S)+c_{+}(S)}{2}\right]^{t-1},\end{split}$ (104)
and finally
$\operatorname{\mathcal{F}}^{+}_{t}(\mathcal{G})=\ \frac{1}{|G|}\sum_{S\in G}\left[\frac{f_{+}(S)+c_{+}(S)}{2}\right]^{t-1}.$ (105)
A.3 Lemma for Theorem 5
@@
Appendix B Sampling p-Cliffords
@@
    We can now assume that $v$ is not fully contained in $w$, or vice versa. If that were the case, they would have a common zero since neither of them can be equal to $j$. Therefore the only other option is that $v_{i}=1,w_{i}=0$ for (at least) one index $i$, and $v_{i^{\prime}}=0,w_{i^{\prime}}=1$ for another index $i^{\prime}$. The vector $a$ that only has ones at $i$ and $i^{\prime}$, and zeroes everywhere else, therefore describes a valid reflection. The other one is again described by $b=w-v-a$.
The actual routine for sampling orthogonal matrices is now straightforward.
@@
It is $\mathcal{O}(N^{3})$ because it has $N$ recursive steps, during each of which it takes $\mathcal{O}(N)$ to find the Householder reflection that maps $(1,0,\ldots,0)^{T}\in\mathbb{F}_{2}^{N}$ to the (odd-parity) binary expansion corresponding to the leading $N-1$ bits of $1\leq i\leq|\operatorname{O}(2n)|$. Applying these reflections to said basis vector (and the other $\mathcal{O}(N)$ ones from the previous recursion steps) then takes $\mathcal{O}(N)$ steps each, since doing so only involves computing inner products and adding vectors. Overall it therefore requires $\mathcal{O}(N^{2})$ time to apply the Householders to all columns. Uniqueness of this mapping follows from the ability to perform all these steps in reverse, i.e. by reading the binary representation of the corresponding index $i$ off the the leading column vectors at each inverse recursion step.

### v3

@@ D.2 Theorem 5
where the second line is a direct consequence of the discrete Fourier transform identity [23]
$\sum_{v\in M}(-1)^{a^{T}v}=|M|\,\delta_{a\in M^{\perp}},$ (127)
which holds for any subspace $M\subseteq\mathbb{F}_{2}^{2n}$ and the corresponding orthogonal space.
$M^{\perp}\coloneqq\{v\in\mathbb{F}_{2}^{2n}\,|\,v^{T}v^{\prime}=0\quad\forall v^{\prime}\in M\}.$ (128)
Equation (127) also implies that
$\begin{split}&\sum_{a\in\mathbb{F}_{2}^{2n,+}}\lvert\operatorname{tr}[\Pi_{+}\operatorname{\Gamma}(S)\operatorname{\mu}(a)]\rvert^{2}\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}\sum_{a\in\mathbb{F}_{2}^{2n,+}}\frac{(-1)^{a^{T}v}}{2}(\delta_{Sv,v}+\delta_{Sv,v^{c}})\\ ={}&\lvert\mathbb{F}_{2}^{2n,+}\rvert\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{1}{2}(\delta_{Sv,v}+\delta_{Sv,v^{c}})\,\delta_{v\in(\mathbb{F}_{2}^{2n,+})^{\perp}}\\ ={}&\lvert\mathbb{F}_{2}^{2n,+}\rvert,\end{split}$ (129)
since $(\mathbb{F}_{2}^{2n,+})^{\perp}\cap\mathbb{F}_{2}^{2n,+}=\{0,j\}$, and both elements are fixed points. Hence the sum in the beginning of (LABEL:eq:sum_trace_squared_parity) must have $2\cdot\lvert\mathbb{F}_{2}^{2n,+}\rvert/(f_{+}(S)+c_{+}(S))$ non-zero terms, each of which is equal to (126). It follows that
$\begin{split}&\sum_{a\in\mathbb{F}_{2}^{2n,+}}\lvert\operatorname{tr}[\Pi_{+}\operatorname{\Gamma}(S)\operatorname{\mu}(a)]\rvert^{2t}\\ ={}&\frac{2\,\lvert\mathbb{F}_{2}^{2n,+}\rvert}{(f_{+}(S)+c_{+}(S))}\cdot\left[\frac{f_{+}(S)+c_{+}(S)}{2}\right]^{t}\\ ={}&\lvert\mathbb{F}_{2}^{2n,+}\rvert\cdot\left[\frac{f_{+}(S)+c_{+}(S)}{2}\right]^{t-1},\end{split}$ (130)
which when plugged into (119) yields
thus completing the proof. ∎
D.3 Lemma for Theorem 6
@@
Appendix E Sampling p-Cliffords
@@
    We can now assume that $v$ is not fully contained in $w$, or vice versa. If that were the case, they would have a common zero since neither of them can be equal to $j$. Therefore the only other option is that $v_{i}=1,w_{i}=0$ for (at least) one index $i$, and $v_{i^{\prime}}=0,w_{i^{\prime}}=1$ for another index $i^{\prime}$. The vector $a$ that only has ones at $i$ and $i^{\prime}$ – and zeroes everywhere else – therefore describes a valid reflection. The other one is again described by $b=w-v-a$.
The actual routine for sampling orthogonal matrices is now straightforward. It is $\mathcal{O}(N^{3})$ because it has $N$ recursive steps, during each of which it takes $\mathcal{O}(N)$ to find the Householder reflection that maps $(1,0,\ldots,0)^{T}\in\mathbb{F}_{2}^{N}$ to the (odd-parity) binary expansion corresponding to the leading $N-1$ bits of $1\leq i\leq|\operatorname{O}(N)|$. Applying these reflections to said basis vector (and the other $\mathcal{O}(N)$ ones from the previous recursion steps) then takes $\mathcal{O}(N)$ steps each, since doing so only involves computing inner products and adding vectors. Overall it therefore requires $\mathcal{O}(N^{2})$ time to apply the Householders to all columns. Uniqueness of this mapping follows from the ability to perform all these steps in reverse, i.e. by reading the binary representation of the corresponding index $i$ off the the leading column vectors at each inverse recursion step.